# Optimising Communication Control Factors for Energy Consumption in Rural V2X

### v2

@@ VI Conclusions
This study considers the balance of safety and energy consumption of rural connected brakes based on optimising three key communication control factors: $\mathrm{SCS}$, ${P}_{\mathrm{t}}$, and $\mathrm{MCS}$. We evaluated $\mathrm{PRR}$ and $D_{\mathrm{comm}}$ under diverse traffic conditions, followed by an energy consumption analysis. The results demonstrate that, an $\mathrm{SCS}$ of $30$ $\mathrm{kHz}$ combined with $\mathrm{MCS}$ 8 consistently ensures reliable $\mathrm{PRR}$ at the safety threshold distance $D_{\mathrm{comm}}$. Under light traffic, where interference is limited, maintaining a lower ${P}_{\mathrm{t}}$ is more energy efficient while still preserving safety margins at $D_{\mathrm{comm}}$. Under heavy traffic, increasing ${P}_{\mathrm{t}}$ from $23$ $\mathrm{dBm}$ to $26$ $\mathrm{dBm}$ can effectively improve safety by increasing $\mathrm{PRR}$ and $D_{\mathrm{comm}}$. However, this improvement comes at the cost of more than a 100100% increase in total energy expenditure to deliver all packets from the RSU to every vehicle, although it provides substantial gains in communication reliability and enables support for higher vehicle densities. These findings underscore the importance of adaptive optimisation of three communication factors to balance the safety and energy consumption of rural connected brakes. Overall, the findings confirm that adaptive configuration of ${P}_{\mathrm{t}}$, $\mathrm{SCS}$, and $\mathrm{MCS}$ is essential to balance energy efficiency with safety-critical performance in rural V2X systems. By explicitly incorporating $D_{\mathrm{comm}}$ as a design metric, this work provides a detail analysis on optimising communication control factors to ensuring a safe, energy-efficient operation of connected braking applications in rural traffic conditions, in order to bring useful insights to the future rural CAV developments.

### v3

@@ VI Conclusions
This study investigates the impact of three key communication control factors: $\mathrm{SCS}$, ${P}_{\mathrm{t}}$, and $\mathrm{MCS}$ on safety-relevant performance metrics and RF energy consumption for rural connected braking. We evaluated $\mathrm{PRR}$ and $D_{\mathrm{comm}}$ under diverse traffic conditions, followed by an energy consumption analysis. The results demonstrate that, an $\mathrm{SCS}$ of $30$ $\mathrm{kHz}$ combined with $\mathrm{MCS}$ 8 consistently ensures reliable $\mathrm{PRR}$ at the safety threshold distance $D_{\mathrm{comm}}$. Under light traffic, where interference is limited, maintaining a lower ${P}_{\mathrm{t}}$ is more energy efficient while still preserving safety margins at $D_{\mathrm{comm}}$. Under heavy traffic, increasing ${P}_{\mathrm{t}}$ from $23$ $\mathrm{dBm}$ to $26$ $\mathrm{dBm}$ can effectively improve safety by increasing $\mathrm{PRR}$ and $D_{\mathrm{comm}}$. However, this improvement comes at the cost of more than a 100100% increase in total energy expenditure to deliver all packets from the RSU to every vehicle, although it provides substantial gains in communication reliability and enables support for higher vehicle densities. These findings underscore the importance of adaptive control of three communication factors to balance the safety and energy consumption of rural connected brakes. Overall, the findings confirm that adaptive configuration of ${P}_{\mathrm{t}}$, $\mathrm{SCS}$, and $\mathrm{MCS}$ is essential to balance energy efficiency with safety-critical performance in rural V2X systems. By explicitly incorporating $D_{\mathrm{comm}}$ as a design metric, this work provides a detail analysis on performance evaluation of communication control factors to ensuring a safe, energy-efficient operation of connected braking applications in rural traffic conditions, in order to bring useful insights to the future rural CAV developments. A future research direction is the development of adaptive configuration mechanisms that dynamically adjust communication control factors according to real-time traffic and channel conditions in rural environments.